# FlashAudio: Rectified Flows for Fast and High-fidelity Text-to-Audio Generation

### v1

@@ Abstract
Recent advancements in latent diffusion models (LDMs) have markedly enhanced text-to-audio generation, yet their iterative sampling processes impose substantial computational demands, limiting practical deployment. While recent methods utilizing consistency-based distillation aim to achieve few-step or single-step inference, their one-step performance is constrained by curved trajectories, preventing them from surpassing traditional diffusion models. In this work, we introduce FlashAudio with rectified flows to learn straight flow for fast simulation. To alleviate the inefficient timesteps allocation and suboptimal distribution of noise, FlashAudio optimizes the time distribution of rectified flow with Bifocal Samplers and proposes immiscible flow to minimize the total distance of data-noise pairs in a batch vias assignment. Furthermore, to address the amplified accumulation error caused by the classifier-free guidance (CFG), we propose Anchored Optimization, which refines the guidance scale by anchoring it to a reference trajectory. Experimental results on text-to-audio generation demonstrate that FlashAudio’s one-step generation performance surpasses the diffusion-based models with hundreds of sampling steps on audio quality and enables a sampling speed of 400x faster than real-time on a single NVIDIA 4090Ti GPU.   [^1] Audio Samples are available at https://FlashAudio-TTA.github.io/.
@@ 1 Introduction
Rectified Flow (RF)(Liu et al., 2022) is a novel flow-based generative model that linearly transfers the source distribution to the target distribution. This property of linear transfer facilitates simulation-free training and faster sampling at inference time. Recently, RF-based methods have garnered significant attention in image synthesis(Liu et al., 2023d; Esser et al., 2024) and text-to-speech applications (Guan et al., 2024). For instance, InstaFlow (Liu et al., 2023d) generates high-quality images in just one step. While this model class offers superior theoretical properties and has proven effective in both image and speech domains, it has not yet been explored for text-to-audio generation.
@@ 3.4.2 Anchored Optimization
Classifier-Free Guidance (CFG) (Ho and Salimans, 2021) plays a critical role in generating audios that align with text prompts. During sampling, Liu et al. (2023d) introduce a guided velocity field in the context of text-conditioned rectified flows using the formula:
@@ A.1 Text-to-audio Generation
Text-to-Audio Generation is an emerging task that has witnessed notable advancements in recent years. For instance, Diffsound Yang et al. (2023) leverages a pre-trained VQ-VAE Van Den Oord et al. (2017) on mel-spectrograms to encode audio into discrete codes, subsequently utilized by a diffusion model for audio synthesis. AudioGen Kreuk et al. (2023b) frames text-to-audio generation as a conditional language modeling task, while Make-An-Audio Huang et al. (2023c), AudioLDM 2 Liu et al. (2023a), and TANGO Ghosal et al. (2023) are all founded on the Latent Diffusion Model (LDM), which significantly enhances sample quality. However, a notable drawback of diffusion models (Liu et al., 2023c, 2024b) lies in their iterative sampling process, leading to slow inference and restricting real-world applications. Recent work has thus focused on consistency distillation to achieve few-step and one-step generation. For instance, AudioLCM leverages consistency distillation with a multi-step ODE solver, achieving two-step generation performance comparable to diffusion models. Other works adopt one-step generation but struggle to align with LDMs. In this work, we introduce a novel rectified flow model for fast and high-fidelity text-to-audio generation and achieve superior performance in both few-step and one-step generation.
@@ A.2 Flow-based Models
Flow matching (Lipman et al., 2022) models the vector field of the transport probability path from noise to data samples. Compared to score-based models like DDPM (Ho et al., 2020b), flow matching offers more stable and robust training, along with superior performance. Specifically, rectified flow matching (Liu et al., 2022) learns the transport ODE to follow straight paths between noise and data points as closely as possible, reducing transport costs and enabling fewer sampling steps through the reflow technique. This approach has shown exceptional performance in accelerating image generation (Liu et al., 2023d; Liu, 2022; Esser et al., 2024). In the realm of audio generation, Voicebox (Le et al., 2024) employs flow matching to develop a large-scale multi-task speech generation model. Its successor, Audiobox (Vyas et al., 2023), expands this approach into a unified audio generation model guided by natural language prompts. VoiceFlow (Guo et al., 2024) introduces rectified flow matching into text-to-speech (TTS), achieving speech generation with fewer inference steps. However, in the task of text-to-audio (TTA) generation, rectified flow models have not yet been explored to enhance sample efficiency and high-quality audio generation.
@@ Appendix C Additional Quantitative Comparison
Due to the page limit, we provide additional quantitative results with performance on text-to-music generation, scalable model size of transformer backbone, and more ablation results on bifocal samplers.

### v2

@@ Abstract
Recent advancements in latent diffusion models (LDMs) have markedly enhanced text-to-audio generation, yet their iterative sampling processes impose substantial computational demands, limiting practical deployment. While recent methods utilizing consistency-based distillation aim to achieve few-step or single-step inference, their one-step performance is constrained by curved trajectories, preventing them from surpassing traditional diffusion models. In this work, we introduce FlashAudio with rectified flows to learn straight flow for fast simulation. To alleviate the inefficient timesteps allocation and suboptimal distribution of noise, FlashAudio optimizes the time distribution of rectified flow with Bifocal Samplers and proposes immiscible flow to minimize the total distance of data-noise pairs in a batch vias assignment. Furthermore, to address the amplified accumulation error caused by the classifier-free guidance (CFG), we propose Anchored Optimization, which refines the guidance scale by anchoring it to a reference trajectory. Experimental results on text-to-audio generation demonstrate that FlashAudio’s one-step generation performance surpasses the diffusion-based models with hundreds of sampling steps on audio quality and enables a sampling speed of 400x faster than real-time on a single NVIDIA 4090Ti GPU. Code will be available at https://github.com/liuhuadai/FlashAudio.   [^1] Audio Samples are available at https://FlashAudio-TTA.github.io/.
@@ 1 Introduction
Rectified Flow (RF)(Liu et al., 2022) is a novel flow-based generative model that linearly transfers the source distribution to the target distribution. This property of linear transfer facilitates simulation-free training and faster sampling at inference time. Recently, RF-based methods have garnered significant attention in image synthesis(Liu et al., 2023e; Esser et al., 2024) and text-to-speech applications (Guan et al., 2024). For instance, InstaFlow (Liu et al., 2023e) generates high-quality images in just one step. While this model class offers superior theoretical properties and has proven effective in both image and speech domains, it has not yet been explored for text-to-audio generation.
@@ 3.4.2 Anchored Optimization
Classifier-Free Guidance (CFG) (Ho and Salimans, 2021) plays a critical role in generating audios that align with text prompts. During sampling, Liu et al. (2023e) introduce a guided velocity field in the context of text-conditioned rectified flows using the formula:
@@ 6 Limitation And Potential Risks
Acknowledgements
The research was supported by NSFC (No.62206234) from Mainland China, Early Career Scheme (ECS-HKUST22201322) from Hong Kong RGC, and National Natural Science Foundation of China under Grant No.62222211 and No.U24A20326.
@@ A.1 Text-to-audio Generation
Text-to-Audio Generation is an emerging task that has witnessed notable advancements in recent years. For instance, Diffsound Yang et al. (2023) leverages a pre-trained VQ-VAE Van Den Oord et al. (2017) on mel-spectrograms to encode audio into discrete codes, subsequently utilized by a diffusion model for audio synthesis. AudioGen Kreuk et al. (2023b) frames text-to-audio generation as a conditional language modeling task, while Make-An-Audio Huang et al. (2023c), AudioLDM 2 Liu et al. (2023a), and TANGO Ghosal et al. (2023) are all founded on the Latent Diffusion Model (LDM), which significantly enhances sample quality. However, a notable drawback of diffusion models (Liu et al., 2023d, 2024b) lies in their iterative sampling process, leading to slow inference and restricting real-world applications. Recent work has thus focused on consistency distillation to achieve few-step and one-step generation. For instance, AudioLCM leverages consistency distillation with a multi-step ODE solver, achieving two-step generation performance comparable to diffusion models. Other works adopt one-step generation but struggle to align with LDMs. In this work, we introduce a novel rectified flow model for fast and high-fidelity text-to-audio generation and achieve superior performance in both few-step and one-step generation.
@@ A.2 Flow-based Models
Flow matching (Lipman et al., 2022) models the vector field of the transport probability path from noise to data samples. Compared to score-based models like DDPM (Ho et al., 2020b), flow matching offers more stable and robust training, along with superior performance. Specifically, rectified flow matching (Liu et al., 2022) learns the transport ODE to follow straight paths between noise and data points as closely as possible, reducing transport costs and enabling fewer sampling steps through the reflow technique. This approach has shown exceptional performance in accelerating image generation (Liu et al., 2023e; Liu, 2022; Esser et al., 2024). In the realm of audio generation, Voicebox (Le et al., 2024) employs flow matching to develop a large-scale multi-task speech generation model. Its successor, Audiobox (Vyas et al., 2023; Ye et al., 2024), expands this approach into a unified audio generation model guided by natural language prompts. VoiceFlow (Guo et al., 2024) introduces rectified flow matching into text-to-speech (TTS), achieving speech generation with fewer inference steps. However, in the task of text-to-audio (TTA) generation, rectified flow models have not yet been explored to enhance sample efficiency and high-quality audio generation.
@@ Appendix C Additional Quantitative Comparison
Due to the page limit, we provide additional quantitative results with performance on text-to-music generation, scalable model size of transformer backbone (Vaswani et al., 2023; Liu et al., 2023c), and more ablation results on bifocal samplers.